# Improving the Language Understanding Capabilities of Large Language Models Using Reinforcement Learning

### v2

@@ 5.4 Results on SuperGLUE Benchmark
As shown in Table 2, the PPO-tuned LLAMA2-7B achieved the highest average performance, surpassing all baselines. PPO demonstrated particularly strong improvements on reasoning-intensive tasks like COPA and MultiRC, where it significantly outperformed both prompting methods and encoder-only models. These results highlight the effectiveness of PPO in enhancing the model’s capabilities, particularly for tasks requiring reasoning and contextual understanding.
Another interesting observation is that few-shot prompting has resulted in reduced performance compared to zero-shot prompting on various challenging tasks such as COPA, MultiRC, WiC, and WSC. This suggests that the demonstrations provided alongside the task-specific prompt may not be effectively guiding the model toward the correct answer and could potentially be introducing confusion instead of clarifying how to approach the task.
5.5 Performance Comparison Across Different LLMs
To assess the consistency of our findings across different models, we evaluated Qwen2.5-7B and MPT-7B alongside LLAMA2-7B on the STS-B dataset from the GLUE benchmark and the COPA dataset from the SuperGLUE benchmark. The results confirm that PPO-based fine-tuning consistently outperforms the BERT-large model, as well as the zero-shot and few-shot prompting baselines for all LLMs, highlighting its effectiveness across different LLMs. Notably, another independent observation is that for all LLMs, few-shot prompting underperforms zero-shot prompting on the COPA dataset. This aligns with our earlier findings in section 5.4, suggesting that the few-shot examples may be introducing noise rather than improving performance in these reasoning tasks.
Models STS-B COPA
BERT-large 86.5 70.6
LLAMA2-7B
Zero-shot prompting 27.5 57.0
Few-shot prompting 45.5 47.6
PPO-ST 92.6 88.6
Qwen2.5-7B
Zero-shot prompting 83.7 96.6
Few-shot prompting 87.0 47.6
PPO-ST 92.2 97.0
MPT-7B
Zero-shot prompting 19.7 57.4
Few-shot prompting 21.7 49.6
PPO-ST 89.3 84.0
Table 3: Performance comparison of LLAMA2-7B, Qwen2.5-7B Hui et al. (2024), and MPT-7B Team (2023) models on the GLUE STS-B and SuperGLUE COPA tasks under zero-shot prompting, few-shot prompting, and PPO based fine-tuning. Results are sourced from the official GLUE benchmark and SuperGLUE benchmark evaluation servers. For STS-B, we report Spearman correlation, and for COPA, accuracy is used as the evaluation metric.
6 Conclusion
Prompting-based approaches, such as zero-shot and few-shot prompting, have gained popularity for adapting LLMs to downstream tasks. However, when applied to LLAMA2-7B, these methods underperform on NLU tasks compared to smaller encoder-only models like BERT-base and BERT-large. To address this limitation, we explore two fine-tuning strategies that leverage LoRA layers to reduce computational overhead. First, we employ supervised fine-tuning by concatenating task-specific prompts, input queries, and ground-truth labels, optimizing the model with the next-token prediction objective. While this approach improves LLAMA2-7B’s performance over prompting-based methods, it still lags behind BERT-base on the GLUE benchmark. To further enhance performance, we adopt PPO, treating the LLM as a policy that generates the next token (action) based on the current input sequence (state). A reward function then evaluates how closely the generated tokens match the ground-truth labels, guiding updates to the policy. PPO based fine-tuning of LLAMA2-7B, tested across benchmarks like GLUE, and SuperGLUE, resulted in significant performance gains, outperforming strong baselines like BERT-large. Similar trends were observed in other LLMs, including Qwen2.5-7B and MPT-7B, showcasing the robustness of this approach. These findings underscore the effectiveness of PPO in enhancing NLU capabilities in LLMs. Future work could extend these techniques to more diverse datasets and refine reward functions for handling complex NLU tasks.
7 Reproducibility Statement
We provide our codes at https://anonymous.4open.science/r/LLM_NLU-BE83. In the code repo, we provide instructions on how to reproduce experimental results. Furthermore, we also provided the hyperparameter settings in section 5.1.
Appendix A Reward Curve for PPO Fine-Tuning in a Multitask Setting on the GLUE Dataset
We present the reward curve from fine-tuning LLAMA2-7B using PPO in a multitask setting on the GLUE dataset. Figure 2 illustrates the reward values over training iterations, offering insights into the training dynamics of the model. The curve serves as a key performance metric, tracking the model’s learning progress across multiple tasks. The consistent upward trend demonstrates that PPO fine-tuning effectively improves LLAMA2-7B’s ability to generate task-relevant outputs.
Figure 2: Reward curve for multitask PPO fine-tuning of LLAMA2-7B on the GLUE dataset. The plot illustrates the relationship between training iterations (x-axis) and reward values (y-axis), demonstrating the effectiveness of the PPO optimization approach in enhancing model performance over time.

### v4

@@ Appendix C Results on SuperGLUE Benchmark
As shown in Table 6, the PPO-tuned LLAMA2-7B-chat-hf achieved the highest average performance, surpassing all baselines. PPO demonstrated particularly strong improvements on reasoning-intensive tasks like COPA and MultiRC, where it significantly outperformed both prompting methods and encoder-only models. These results highlight the effectiveness of PPO in improving the model’s capabilities, particularly for tasks requiring reasoning and contextual understanding.
It is worth noting that on MultiRC, few-shot prompting performs slightly worse than zero-shot prompting. This may be because MultiRC involves long input contexts, and incorporating multiple examples in a few-shot prompt can cause the total input length to approach or exceed the LLMs maximum context window. Even in the one-shot setting, providing an excessively long context can dilute the model’s attention, potentially leading to reduced performance.
Appendix D Evaluation on Reading Comprehension Tasks
We evaluate LLAMA2-7B-chat-hf on the SQuAD reading comprehension task, where the objective is to select a passage from a given context that best answers a question. Two training strategies are compared: Supervised Fine-Tuning (SFT), which directly uses the ground-truth answer as the training label, and Proximal Policy Optimization (PPO), which leverages reward functions based on Exact Match (EM) and F1 Score. EM metric is computed by comparing a normalized prediction against the normalized ground truth (with normalization involving lowercasing and punctuation removal); a perfect match yields an EM score of 1, otherwise 0. F1 score measures word-level overlap, balancing how many predicted words are correct (precision) and how many ground-truth words are included (recall).
Models were fine-tuned for one epoch on the SQuAD training set and evaluated on the development set. In our evaluation, zero-shot prompting yields an EM of 7.66 and an F1 score of 32.27. SFT significantly improves these metrics (EM: 59.17, F1: 76.48), while PPO further enhances performance, achieving an EM of 65.74 and an F1 score of 81.82—corresponding to improvements of 6.57 and 5.34 points over SFT, respectively.
These results indicate that optimizing with reward functions based on EM and F1 via PPO leads to further improvements in reading comprehension performance, thereby validating our approach relative to both zero-shot prompting and standard SFT.
Method EM F1
Original 7.66 32.27
SFT 59.17 76.48
PPO 65.74 81.82
Table 7: Performance of LLAMA2-7B-chat-hf on the SQuAD dataset. PPO uses Exact Match and F1 as reward signals. Best results are highlighted in bold.
Appendix E Impact of Fine-Tuning on Language Modeling Ability
We experiment with SFT and PPO to improve NLU capabilities of LLMs and observe improved performance using PPO. However, it is crucial to ensure that fine-tuning methods do not significantly degrade the models’ general language generation abilities. To assess this, we directly evaluate the PPL (jel, 1977; Chelba and Jelinek, 2000) of LLAMA2-7B-chat-hf models fine-tuned on the SST-2 dataset using the WikiText-2 test set (Merity et al., 2016), which follows a natural human-written text distribution. We compare these fine-tuned models against the original, non-fine-tuned baseline model, with the expectation that the PPL of the fine-tuned models should closely match the baseline. Our results reveal that the original LLAMA2-7B-chat-hf achieves a perplexity of 6.939. The PPO-fine-tuned model closely maintains this baseline performance with a perplexity of 6.966, indicating minimal impact on its general language modeling capabilities. In contrast, the SFT-fine-tuned model displays a notably higher perplexity of 7.384, suggesting a significant reduction in generation capabilities due to convergence toward task-specific training distributions. We conjecture that PPO’s clipping mechanism effectively constrains policy updates, preventing large deviations from the reference model and thereby preserving the original language modeling capabilities of LLMs. These findings underscore PPO’s effectiveness in maintaining the general language abilities of LLMs during fine-tuning.
Method perplexity
Original 6.939
SFT 7.384
PPO 6.966
Table 8: Perplexity of LLAMA2-7B-chat-hf on the WikiText-2 test set. Lower perplexity indicates better language modeling ability.
Appendix F Comparison of RL Algorithms: PPO vs. GRPO
Algorithm SST-2 MRPC (F1) RTE CoLA QNLI Avg. Per-Step Runtime (s)
SFT 73.8 85.8 80.4 50.7 93.6 76.9 4.124
PPO 96.4 89.4 84.3 59.9 93.2 84.6 4.299
GRPO 96.7 91.2 88.5 55.2 93.1 84.9 5.155
Table 9: Zero-shot GLUE performance (SST-2, MRPC F1, RTE, CoLA, QNLI) and per-step runtime. Best score in each column is bolded.
Our objective is to improve the natural language understanding capabilities of the base (policy) model through RL fine-tuning. In this context, we compare two approaches: PPO and Group Relative Policy Optimization (GRPO) (Shao et al., 2024). PPO is highly effective but introduces additional computational overhead. This overhead stems from the need for repeated sampling and from updating a separate critic model to compute value functions. In contrast, GRPO was designed to mitigate these costs by bypassing the critic model entirely. Instead of maintaining a separate value network, GRPO samples multiple trajectories per prompt and computes each trajectory’s advantage by comparing its reward to the batch’s average (and standard deviation). This method not only simplifies the architecture but also reduces memory usage.
For our experiments, we utilized the TRL library (von Werra et al., 2020) on a single Nvidia A100 GPU, with a batch size of 16 and gradient checkpointing enabled. While SFT involves a simple forward pass, loss computation, and backward pass per step, both PPO and GRPO add extra steps such as LLM sampling, reward calculation, and advantage estimation.
As detailed in Table 9, both PPO and GRPO deliver notable performance improvements over SFT. Notably, PPO only incurs about a 4% increase in per-step runtime compared to SFT. However, GRPO’s need to generate multiple responses per sample results in a higher runtime, despite its memory efficiency benefits. Overall, our analysis highlights the trade-offs between these RL algorithms: PPO offers efficient runtime with the cost of additional overhead from the critic model, while GRPO reduces memory usage at the expense of increased sampling time.
Appendix G Reward Function Design and Evaluation
Method Accuracy (%)
PPO 96.4
PPO-RM 89.7
(a) SST-2 performance on GLUE.
Method GPT Eval. Score
PPO 3.479
PPO-RM 4.104
(b) Quality of generated analyses.
Table 10: Comparison of reward function designs for LLAMA2-7B-chat-hf. The model trained with a rule based reward (PPO) achieves a high SST-2 classification accuracy of 96.4%, while incorporating a sophisticated reward model (PPO-RM) significantly reduces accuracy (89.7%) but yields substantially improved analysis quality, with a GPT evaluation score of 4.104 compared to 3.479 for the simple reward. Best results are highlighted in bold.
While our primary reward function is based on matching generated outputs to true labels, we recognize that more sophisticated reward designs may be necessary for complex NLU tasks. To address this, we investigate the effect of integrating a reward model into our PPO training, with the aim of enhancing not only classification performance on SST-2 but also the quality of generated analyses.
Reward Modeling Setup. For the first 5,000 training samples of the SST-2 dataset, LLAMA2-7B-chat-hf generates four responses per data point. Each response includes a sentiment judgment (Positive/Negative) and a supporting analysis. To robustly rank these responses, we use GPT-4o as an evaluator. GPT-4o ranks the responses based on: (i) the correctness of the sentiment judgment (i.e., matching the ground truth), (ii) the consistency between the judgment and its accompanying analysis, and (iii) the overall factual correctness and helpfulness of the analysis. To ensure clear differentiation, we include two reference responses—one with only the correct answer and one with only the incorrect answer—and define the ranking order as: correct answer with analysis > only correct answer > incorrect answer with analysis > only incorrect answer.
Training the Reward Model. A reward model is then trained on this ranked dataset using a BERT-based architecture (bert-base-cased). For each input $x$, we consider pairs of responses $(y_{w},y_{l})$, where $y_{w}$ denotes a response ranked higher by our evaluator (GPT-4o) due to its correct sentiment and coherent analysis, and $y_{l}$ denotes a lower-ranked response. The model learns to assign higher scores to better responses via a pairwise ranking loss:
$\begin{split}L(\theta)&=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\\ &\quad\Bigl{[}\log\sigma\bigl{(}r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})\bigr{)}\Bigr{]},\end{split}$ (3)
where $r_{\theta}(x,y)$ is the score assigned to response $y$ given $x$, and $\sigma$ is the sigmoid function converting the score difference into a probability. This loss encourages the reward model to output higher scores for responses with superior judgments and analyses.
Incorporating the Reward Model into PPO Training. During PPO training on SST-2, LLM is tasked with generating both a sentiment judgment and an analysis. The trained reward model provides the reward signal by scoring these outputs. As shown in Table 10(a), while the PPO model trained with reward signals from the reward model (PPO-RM) produces analyses of higher quality, it suffers from a significant reduction in classification performance, dropping from 96.4% to 89.7%. We believe this discrepancy might be due to the limited sample size used for reward model training and potential reward hacking Amodei et al. (2016) during optimization. However, we will explore this further in our future works.
Evaluation of Generated Analyses. To further assess the impact of our reward design, we evaluated the quality of generated analyses. We sampled 100 data points from three models: the original LLAMA2-7B-chat-hf, the PPO model trained using only correct-answer rewards (PPO), and the PPO model trained with the reward model (PPO-RM). GPT-4o then scored each analysis on a scale from 1 to 5 based on answer correctness and logical coherence. As indicated in Table 10(b), the PPO model using reward model signals achieved the highest average score, suggesting that a more complex reward function can enhance the quality of generated outputs.
In summary, while the integration of a reward model in PPO training significantly reduces classification performance compared to using only correct-answer rewards, it considerably improves the GPT evaluation scores of the analyses produced by the LLM.